OpenStudy (anonymous):

Would you make more interest by investing two thousand five hundred dollars in a 6-month certificate of deposit (CD) at three percent interest or by investing one thousand dollars in a twelve-month CD at four percent interest? Show your work.

OpenStudy (anonymous):

I think that the 12 month one will get you more interest. I am not sure what CD stands for. But how I go this: (2500*0.003)*6=450 (1000*0.04)*12=480 Is this correct????